OpenStudy (anonymous):

Which type of triangle has side lengths of 2 in, 5 in, and 6 in?

OpenStudy (lukecrayonz):

 2^2+5^2=6^2 4+25=36 29=/=36 So it's not a right. lol

OpenStudy (lukecrayonz):

Thats all I got.

OpenStudy (p0sitr0n):

we assume 6 is the longest side, so maybe hypotenuse as we have 3 different sides, so not equilateral nor isocel must be a triangle with 3 different sides, where 1 angle > 90, and 2 others <90

OpenStudy (anonymous):

 OBTUSE